OpenStudy (elisaneedshelp):

@OrthodoxMan The lengths of the three sides of a right triangle are consecutive integers. Find them. Idk how to do this, do you know?

OpenStudy (wolf1728):

I happen to know that a right triangle of sides 3, 4 and 5 is one that is used quite often. It just so happens that it has consecutive integers. I have no idea how you would solve this in any other way besides knowing about the 3, 4, 5 triangle.

OpenStudy (elisaneedshelp):

This is such a stupid question that i dont need for the rest of my life

OpenStudy (elisaneedshelp):

hey @annie.kopliku do you know this question?

OpenStudy (anonymous):

Yes I actually do, but it is so complicated to understand. My teacher showed me how to do this, and I had to go over it like 20 times to actually comprehend it. Hold on while I type this out.

OpenStudy (anonymous):

Yes I know how to do this. No I don't know what took me so long to find I had a dangling notification.

OpenStudy (anonymous):

So you know that consecutive numbers are numbers that come after each other. Such as 1,2,3,4... Now we'll be using x (because x is used for basically everything; in this case we'll let x be 0) Do you notice how x=0 x+1= 1 x+2= 2 x+3=3 Those are consecutive numbers. Since the problem says we're dealing with a right triangle, it tells us that a=x b= x+1 c= x+2 Here comes the Pythagorean Theorem.

OpenStudy (anonymous):



OpenStudy (anonymous):

Yeah now you do the pythagorean theorem and solve for x after expanding the equation. Pretty much another polynomial question. YAY!!

OpenStudy (anonymous):

\[x^2+(x+1)2=(x+2)^2\]

OpenStudy (elisaneedshelp):

So interger and number are the same thing just to know? and um so i dont change any number in that equation

OpenStudy (anonymous):

You don't substitute anything and I'm sorry it's supposed to be (x+1)^2 not times 2.

OpenStudy (anonymous):

No. Every integer is a number but every number is not an integer.

OpenStudy (anonymous):

^ How you put that is really confusing.

OpenStudy (elisaneedshelp):

Lol he sounded wise tho xD

OpenStudy (elisaneedshelp):

so am i supposed to try and get x alone with all those x's?

OpenStudy (anonymous):

Do you know how to foil? If not, I'll help you with that. I don't understand why I can't respond immediately, like it won't let me erase my last message so I have to exit this, and wait for you to respond after everything I say.

OpenStudy (anonymous):

Sorry, I have discrete mathematics. My professor talks like that a lot.

OpenStudy (anonymous):

@annie.kopliku That happens when you're not connected to this site properly. I had to refresh my browser to make it work properly.

OpenStudy (elisaneedshelp):

Plz help me with whatever foiling is

OpenStudy (anonymous):

What is foiling?

OpenStudy (anonymous):

All you have to do is expand and solve for x.

OpenStudy (anonymous):

Foiling is like getting that equation and basically "expanding it" you have to put everything on one side, if my OS wasn't malfunctioning I'd be able to show you.

OpenStudy (elisaneedshelp):

:(

OpenStudy (anonymous):

I guess I'll do the foiling.

OpenStudy (anonymous):

I already got it.

OpenStudy (anonymous):

ITS WORKING. YAYAY

OpenStudy (anonymous):

Then showwww meeeeeeeee.

OpenStudy (anonymous):

pweaaaaaaaaaaaaseeeeeeeeeeeee

OpenStudy (anonymous):

\[x^2+(x+1)^2=(x+2)^2\] this turns into \[x^2+(x+1)+(x+1)=(x+2)+(x+2)\]

OpenStudy (elisaneedshelp):

and then what else does it turn in to?

OpenStudy (anonymous):

Sorry, the x^2 turns into x+x Let me redo the foiling. (This is complicated) x + x + x + 1 + x + 1 = x +2 + x + 2 Combine all the like terms

OpenStudy (anonymous):

Hold on let me check this again. I need to do this on paper. I'm so sorry.

OpenStudy (anonymous):

\[x^2+2x+2=2x+4\] \[x^2 +2x - 2x +2 - 4\]

OpenStudy (anonymous):

I know the answer is 3,4,5 that's usually how it always is. But I'm figuring out the steps.

OpenStudy (anonymous):

Yeah I tried that.

OpenStudy (anonymous):

\[x^2 - 2 = 0\]

OpenStudy (anonymous):

Well I think i see why that doesnt work...

OpenStudy (anonymous):

OH WAIT. You did the quadratic thing wrong

OpenStudy (anonymous):

Just know that the answer is 3 4 5 How you get there is gonna be tricky.

OpenStudy (elisaneedshelp):

I can tell its super tricky :/

OpenStudy (anonymous):

\[x^2 + (x+1)^2 = (x+2)^2\] \[x^2 + x^2 + 2x +1 = x^2 + 4x + 4\] \[2x^2 +2x + 1 = x^2 + 4x +4\] \[2x^2 - x^2 +2x - 4x + 1 - 4 = 0\] \[x^2 - 2x - 3 = 0\]

OpenStudy (anonymous):

\[a^2+b^2=c^2 \] is what I started off with. Then I did, \[(x^2) + (x+1)^2 = (x+2)^2\] I "expanded" and got \[(x+x) + (x+x+2) = (x+x+4)\] Combining all the like terms, I got \[4x + 2 = 2x + 4\] Simplifying it I get, \[2x = 2 \] Dividing by 2 I get \[x=1\] Do you think I should substitute and get the other answers @OrthodoxMan

OpenStudy (elisaneedshelp):

Omg how does my teacher give me this if she didnt teach it >_<

OpenStudy (anonymous):

Alot of teachers do this later on, what grade are you in? And what math do you take?

OpenStudy (anonymous):

Actually that isn't right. I think it would work though let me look at it again.

OpenStudy (elisaneedshelp):

Im in 10th grade geometry

OpenStudy (anonymous):

I seee.

OpenStudy (anonymous):

Yeah it isn't right. You didn't deal with your quadratics right. Remember. x*x = x^2 and x+x = 2x.

OpenStudy (anonymous):

Go back to my answer. \[x^2 -2x -3 = 0\]

OpenStudy (anonymous):

I GOT THIS NOW YOU FOIL.

OpenStudy (anonymous):

my teacher taught us this (x-3) (x+1) when you come upon those types of equations.

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

\[3^2+(3+1)^2=(3+2)^2\]

OpenStudy (anonymous):

Factoring that we get: \[x^2 +x -3x -3\] \[(x^2 +x) + (-3x - 3)\] \[x(x+1)-3(x+1)\] \[(x-3)(x+1)\]

OpenStudy (anonymous):

yeah that's what I got on my paper.

OpenStudy (anonymous):

x = -1 is not applicable. So the answer is x = 3.

OpenStudy (anonymous):

Teamwork.

OpenStudy (anonymous):

HOORAY FOR TEAMWORK.

OpenStudy (anonymous):

*high fives*

OpenStudy (anonymous):

From here do your consecutives.

OpenStudy (elisaneedshelp):

omg lol now i have to try and process this in my poor brain

OpenStudy (anonymous):

*Raises hand to defend face from being high-fived*

OpenStudy (anonymous):

If you need help processing. You know who to ask. The team of two just for you.....that was really cheesy.

OpenStudy (anonymous):

That also rhymed, so..

OpenStudy (elisaneedshelp):

Lol yeah you do make a good team

OpenStudy (anonymous):

Yeah then I guess we're pretty good then. Sorry if my absence scared you guys. BATTERY DYING :O:O